OpenStudy (gavin39):

OK I seriously need help on this. Ill do absolutely anything!!!! 1.A mysterious box is delivered to the dinner party you are attending. The label on the box says that the volume of a box is the function f(x) = x3 + 3x2 – 10x – 24. To open the box, you need to identify the correct factors of f(x). Party-goers offer up solutions, and it is your job to find the right ones. Their suggestions are: •(x – 1) •(x + 2) •(x – 3) •(x + 4) •(x + 6) •(x – 12) List the correct factors. Then justify your selections with complete sentences.

OpenStudy (gavin39):

Ill give medals and everything!

OpenStudy (gavin39):

Also need questions 2, 3 , and 4 done. Ill pay!

OpenStudy (gavin39):

2.Three party-goers are in the corner of the ballroom having an intense argument. You walk over to settle the debate. They are discussing a function g(x). You take out your notepad and jot down their statements. •Professor McCoy: She says that 2 is a zero of g(x) because long division with (x + 2) results in a remainder of 0. •Ms. Guerra: She says that 2 is a zero of g(x) because g(2) = 0. •Mr. Romano: He says that 2 is a zero of g(x) because synthetic division with 2 results in a remainder of 0.

OpenStudy (gavin39):

3.Dr. Collier summons you over to his table. He wants to demonstrate the graph of a fourth-degree polynomial function, but the batteries in his graphing calculator have run out of juice. Explain to Dr. Collier how to create a rough sketch of a graph of a fourth-degree polynomial function.

OpenStudy (gavin39):

4.Mrs. Collins is at the table with you and states that the fourth-degree graphs she has seen have four real zeros. She asks you if it is possible to create a fourth-degree polynomial with only two real zeros. Demonstrate how to do this and explain your steps.

OpenStudy (gavin39):

@Directrix

OpenStudy (anonymous):

hey for # 1 i can do it but idk if i can explain it if thats alright

OpenStudy (gavin39):

Yeah that's fine. I was looking for someone to do since I have no idea how or what it is @alec.mccarty

OpenStudy (anonymous):

well you have to use there of the options multiplied together to get that polynomial

OpenStudy (anonymous):

three***

OpenStudy (gavin39):

But I have to do all 6 of them

OpenStudy (anonymous):

no you need to pick the right 3 which when multiplied together give you the polynomial in the queation

OpenStudy (anonymous):

here let me do it then ill let you work it out

OpenStudy (gavin39):

OOOO ok if that's the case then which 3 is it lol

OpenStudy (anonymous):

not sure yet 1 min

OpenStudy (gavin39):

ok

OpenStudy (anonymous):

not x-12

OpenStudy (anonymous):

•(x + 2) •(x – 3) •(x + 4)

OpenStudy (anonymous):

those 3

OpenStudy (gavin39):

OK cool and for the couple complete sentences how should I say I got my answer

OpenStudy (gavin39):

Or how I justify my answers

OpenStudy (anonymous):

I picked 3 of there suggestions because to get to x cubed i knew i would have to have three factors. Also i knew that the numbers in the factors would have to multiply to 24 because of the polynomial. The 24 was negative which meant i would need either 3 negative numbers or one negative number. This ruled out everything but the factor which were correct.

OpenStudy (anonymous):

if u need more i can go more in detail. i just kinda wrapped it up bc it was getting long

OpenStudy (gavin39):

Think you could help me with a couple more of these problems @alec.mccarty

OpenStudy (anonymous):

let me look at em

OpenStudy (gavin39):

Please do man I really need this done or im screwed in class

OpenStudy (anonymous):

i think th 2nd one is ms. G and i dont remember forth degree sorry

OpenStudy (anonymous):

let me think for a couple mins

OpenStudy (gavin39):

Ok no prob

OpenStudy (anonymous):

ok #3

OpenStudy (gavin39):

Sweet!! Find it?

OpenStudy (anonymous):

First off he need to use the leading coefficient test to determine if the side go up or down. If a<0 they go down and if a>0 they go up. Then he needs to factor the polynomial to determine the zeros of the graph. After he factors he needs to find the multiplicity of each of the zeros which can be done by simply looking at the power each factor is raised to. After he plots his point then he needs to pick a point between each of the zeros and plug it into the function to get points on each of the curves. Then he just needs to connect the dots.

OpenStudy (anonymous):

and now #4

OpenStudy (gavin39):

Bro your the best!

OpenStudy (anonymous):

Yes it is possible to only have two real zeros if you have two zeros where the line touches the x axis but does not cross. This is achieved if when you factor the function you end up with two factors which each are squared. When the factors are squared it means that their multiplicity is 2. This means that each of these points count as 2 zeros.

OpenStudy (anonymous):

that should do

OpenStudy (anonymous):

and idk if 2 was right

OpenStudy (gavin39):

Yeah for number to I need to explain it which if you don't know how to no way in hell I will lol but bro your seriesly are a life saver. Really appreciate it

OpenStudy (gavin39):

Now I have one more assignment that looks easy to do but since I don't know any of this I obviously don't know how to do it. If your up for it

OpenStudy (anonymous):

i can explain 2 lol 1 min

OpenStudy (gavin39):

OK cool that would help a lot!

OpenStudy (anonymous):

Ms. G is correct because she demonstrated that when 2 was plugged into the equation the y value is 0 which would make it a zero of the function

OpenStudy (anonymous):

idk about it but it may get you a point or 2

OpenStudy (gavin39):

Fine with me man. Up for the one question on my other assignement?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

ima go smoke a cig real fast ill be back in 10min or so

OpenStudy (gavin39):

ok cool!

OpenStudy (gavin39):

You must: •Label and display your new polynomial identity •Prove that it is true through an algebraic proof, identifying each step •Demonstrate that your polynomial identity works on numerical relationships WARNING! No identities used in the lesson may be submitted. Create your own. See what happens when different binomials or trinomials are combined. Below is a list of some sample factors you may use to help develop your own identity. •(x – y) •(x + y) •(y + x) •(y – x) •(x + a) •(y + b) •(x2 + 2xy + y2) •(x2 – 2xy + y2) •(ax + b) •(cy + d)

OpenStudy (anonymous):

ok im back

OpenStudy (gavin39):

Cool I the question is above your comment

OpenStudy (anonymous):

do you know how complicated it has to be?

OpenStudy (gavin39):

Nope. That is what it all said for the question

OpenStudy (gavin39):

here is another part: You are going to design an advertisement for a new polynomial identity that you are going to invent. Your goal for this activity is to demonstrate the proof of your polynomial identity through an algebraic proof and a numerical proof in an engaging way! Make it so the whole world wants to purchase your polynomial identity and can't imagine living without it! You may do this by making a flier, a newspaper or magazine advertisement, making an infomercial video or audio recording, or designing a visual presentation for investors through a flowchart or PowerPoint.

OpenStudy (anonymous):

thats stupid as pellet but ok give me a few mins

OpenStudy (anonymous):

pellet****

OpenStudy (gavin39):

OK sounds good

OpenStudy (anonymous):

do you have a list of the identities that were in the lesson?

OpenStudy (gavin39):

Unfortunately I don't :/

OpenStudy (anonymous):

ill just make a weird 1 then

OpenStudy (gavin39):

lol ha that's fine

OpenStudy (anonymous):

im not quite sure if this is what it wants but here

OpenStudy (gavin39):

Hey its worth a shot

OpenStudy (anonymous):

Identity y^2+x^2+2xy=(x+y)^2 Algebraic Proof y^2+x^2+2xy=(x+y)^2 Simplify right side y^2+x^2+2xy=(x+y)(x+y) Foil right side y^2+x^2+2xy=xy+y^2+x^2+xy Simplify like terms y^2+x^2+2xy=y^2+x^2+2xy Numerical proof y^2+x^2+2xy=(x+y)^2 (where x= and y=2) 2^2+3^2+2*3*2=(3+2)^2 4+9+12=5^2 25=25

OpenStudy (gavin39):

Bro your the best!!!

OpenStudy (anonymous):

thanks! anything else?

OpenStudy (anonymous):

where x=3 btw

OpenStudy (gavin39):

Hmm there is one more assignment but I wuna give u a break for the night lol

OpenStudy (gavin39):

And x=3?

OpenStudy (anonymous):

for the numerical proof i put "where x= and y=2"

OpenStudy (gavin39):

ooo ok ill fix that

OpenStudy (anonymous):

alright well if your done for tonight ill see you later

OpenStudy (gavin39):

Ok cool. Im going to need help on one more assignment for tomorrow. Cool?

OpenStudy (anonymous):

what time? what time zone are you in?

OpenStudy (gavin39):

Im in florida so I think eastern and not 100% on time sine I have practice tomorrow What times good for you? Im usually up till 3 am

OpenStudy (anonymous):

im in central ill try to get on at 10 or 11so 11-12 for you

OpenStudy (gavin39):

OK ill post the questions up and tag u in them. Cool?

OpenStudy (anonymous):

alright, this is my first time using this site will it send me like a message or sumthin?

OpenStudy (gavin39):

It'll send u a notification. Youll get an email saying u have a notification and or if your on on the top left corner what looks like a sweeping pan. That will have a little one on it to show u have one

OpenStudy (anonymous):

alright sounds good cya then

OpenStudy (gavin39):

Have a goodnight bro

OpenStudy (anonymous):

it is possible to have four real zeroes??

OpenStudy (anonymous):

on fourth degree graohs

OpenStudy (anonymous):

what was the answer?

OpenStudy (anonymous):

The information included in this question is the intellectual property of Florida Virtual School. Posting questions from any FLVS course and receiving answers is considered cheating.